OpenStudy (anonymous):

1. Part 1: Solve the inequality -3x + 4 <= x + 8 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. 2. Part 1: Solve the compound inequality 3x + 6 > 36 and 3x + 6 <= 54 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1

OpenStudy (anonymous):

3. Part 1: Use complete sentences to describe a real-world scenario that could be represented by the inequality 5x + 2y >= 45. Part 2: Choose one ordered pair that is a solution to the given inequality and explain what that ordered pair means in the context of your real-world scenario.

OpenStudy (anonymous):

-3x+4<=x+8 -3x-x<=8-4 -4x<=4 x>=-1

OpenStudy (anonymous):

For the part 2 you dont have to worry bout that i can do it i just need the part 1's

OpenStudy (anonymous):

3x + 6 > 36 and 3x + 6 <= 54 3x>36-6 3x>30 x>10 3x + 6 <= 54 3x<=54-6 3x<=48 x<=16 10<x<=16

OpenStudy (anonymous):

is that gonna be open circle on 10, closed circle on 16 and shading in between?

OpenStudy (anonymous):